OpenStudy (anonymous):

5. –2(x – 3) ≥ 5 – (x + 3) 7. Describe a real life scenario where inequalities are used and state the inequality.

hero (hero):

5 - (x + 3) ≤ -2(x - 3) 5 - x - 3 ≤ -2x + 6 5 - 3 - x ≤ -2x + 6 2 - x ≤ -2x + 6 2x - x ≤ 6 - 2 x ≤ 4

hero (hero):

For real life scenarios using inequalities, it is best to think of ranges between two quantities that represent an ideal situation. For example \(37^{\circ} \le x \le 41^{\circ}\) is the ideal temperature range for food storage in a refrigerator.

hero (hero):

If food is stored below 37 degrees, then freezing could occur. Above 41 degrees, and spoilage could occur.

hero (hero):

Another is speed limits. On certain highways: \(40 < x < 65\) mph is the ideal speed for safe travel.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Could you explain how you did the inequality this this though: "4. 2(x + 5) > 8x – 8 Distribute the 2 2x + 10 > 8x – 8 Subtract ten from both sides 2x > 8x – 18 Subtract 8x from both sides -6x > -18 Divide by -6 x < 3" Im a little confused on how you did it.

hero (hero):

^I'm pretty sure I did not do that one. I never divide by negative

OpenStudy (anonymous):

I know but could explain on the one you did do.

hero (hero):

If I did it, I would have done this: 8x - 8 < 2x + 10 8x - 2x < 8 + 10 6x < 8 + 10 6x < 18 x < 18/6 x < 3

hero (hero):

Oh, you want me to explain by including the explanations for each step?

OpenStudy (anonymous):

Yes please

hero (hero):

–2(x – 3) ≥ 5 – (x + 3) 1. Express the inequality using a less than equal sign 5 - (x + 3) ≤ -2(x - 3) 2. Distribute -1 across x + 3: 5 - x - 3 ≤ -2x + 6 3. Place like terms next to each other (associate property) 5 - 3 - x ≤ -2x + 6 4. Subtraction Property (5 - 3 = 2) 2 - x ≤ -2x + 6 Add 2x to both sides: Subtract 2 from both sides: 2x - x ≤ 6 - 2 Simplify x ≤ 4

hero (hero):

Using the less than equal sign is a matter of preference, however, it does naturally reflect placement along a number line. For example: x ≤ 4 means x includes 4 and all values to the left of 4: